# Non-uniqueness of phase transitions for graphical representations of the Ising model on tree-like graphs

### v1

@@ 2.2. Generalisations and non-uniqueness of random current phase transitions
Let $F$ be a finite graph and that $v,w$ are two vertices. Let $\{\mu_{x,F}\}_{x\in[0,1]}$ be a family of cut point factorising percolation measures. Suppose that there exists a finite graph $F$ such that $x\mapsto\mu_{x,F}[v\leftrightarrow w]$ is not monotone. Then, there exists an infinite graph $\mathbb{M}$ such that $\mathcal{P}(\mu_{x,\mathbb{M}})$ is disconnected.
@@
For $p\in(0,1)$ let $G_{p}\sim\mathbb{P}_{p,\mathbb{T}^{d}}$. Notice that almost surely, $x_{c}(\mathbb{P}_{x,G_{p}})=\frac{1}{p(d-1)}.$ By tuning the parameters $p$ and $d$ appropriately, we can make sure that
@@
Now, construct $\mathbb{M}$ by sampling $G_{p}$ and substituting each edge of $G_{p}$ by a copy of $F$, gluing in the same way as in our construction in the proof of Theorem 1.1. ∎
@@
As a consequence, we can see that the single random current also admits a disconnected percolation regime: Recall that [19, Figure 2.3] gives an example of a graph for which connection probabilities are not monotone, hence, by the above proposition, it follows that we get:
@@
$\nu_{\beta,G}[\omega]=\frac{1}{Z_{G,\beta}}\beta^{\absolutevalue{\omega}}1\!\!1_{\Omega_{\emptyset}(\omega)=\{0\}},$
where 00 denotes the empty graph, which is even. It is immediate that if $(G_{1},G_{2})$ is a cut-point factorisation of $G$ and $F_{1},F_{2}$ are subforests of $G_{1}$ and $G_{2}$ respectively, then $F_{1}\cup F_{2}$ is a subforest of $G$. Thus, $\nu_{\beta,G}$ cut-point factorises. At present, the connection probabilities of the model are conjectured to be monotone in $\beta$ on all finite graphs $G$ [15, p.2], but supposing that they are not, our construction would go through for this model as well, yielding a graph on which the percolation phase transition of the model is not unique.
@@ 3. Phase transitions of the wired models on the $d$-regular tree coincide
In this section, we prove the first part of Theorem 1.3 on trees. The main tool in the proof is the observation that all wired cycles in $\mathbb{T}^{d}$ are infinite. In particular, $\ell^{1}_{x,\mathbb{T}^{d}}$ is empty if it does not percolate.
@@
By tail-triviality (see [14, Theorem 10.67]), $\phi^{1}_{x,\mathbb{T}^{d}}[\mathcal{C}_{\infty}]\in\{0,1\}$ and since $\phi^{1}_{x,\mathbb{T}^{d}}[\omega_{e}=0]\in(0,1),$ we conclude that
@@
For $x>x_{c}(\phi^{1}_{\mathbb{T}^{d}}),$ there exists $c>0$ such that for any vertex $v$, we have
@@
Consider two neighbouring vertices $v,w$ and define the event $A_{v}=\{v\longleftrightarrow\infty\text{ in }\mathbb{T}^{d}\setminus\{v,w\}\}$ and analogously $A_{w}$. Define the event $L_{v,w}:=A_{v}\cap A_{w}\cap\{(v,w)\text{ open}\}$. In words, this is the event that that there is a loop which contains the edge $(v,w)$ goes through the wired boundary at infinity. Now, suppose $x>x_{c}(\phi^{1}_{\mathbb{T}^{d}})$. In this case, by the FKG inequality, we obtain a lower bound that a random-cluster configuration $\omega$ satisfies $L_{v,w}$:
@@
Second, the probability of a given edge, which is part of a loop in $\omega$, being open in $\operatorname*{UEG}_{\omega}$ is $\frac{1}{2}$ (see the much more general statement [16, Lemma 3.5]).
@@
In order to finish the proof of the first statement in Theorem 1.3, we make a short aside to discuss the subcritical regime of the random-cluster and random current models on the tree. It is classical that, for $x<x_{c}(\phi_{\mathbb{T}^{d}}^{1}),$ we have that $\phi_{x,\mathbb{T}^{d}}^{1}=\mathbb{P}_{x,\mathbb{T}^{d}}$ (see [14, Theorem 10.67]). A similar result holds for the double random current:
@@
Since $x<x_{c}(\phi_{\mathbb{T}^{d}}^{1})$ and $\ell^{1}_{x,\mathbb{T}^{d}}\preceq\phi^{1}_{x,\mathbb{T}^{d}}$, we get that $\ell_{x,\mathbb{T}^{d}}^{1}[\mathcal{C}_{\infty}]\leq\phi_{x,\mathbb{T}^{d}}^{1}[\mathcal{C}_{\infty}]=0.$ Therefore, since all trivial components of an even subgraph of $\mathbb{T}^{d}$ are infinite, we have that
@@
For $d\geq 2$, then
$x_{c}(\ell^{1}_{\mathbb{T}^{d}})=x_{c}(\mathbf{P}^{1}_{\mathbb{T}^{d}})=x_{c}(\mathbf{P}^{1}_{\mathbb{T}^{d}}\cup\mathbf{P}^{1}_{\mathbb{T}^{d}}).$
@@
For $a)$ if $x<x_{c}(\phi^{1}_{\mathtt{C}_{n}^{d}}),$ rather than $\ell^{1}_{x,\mathtt{C}_{n}^{d}}$ being deterministically empty, it includes each simple cycle of $\mathtt{C}_{n}^{d}$ independently since the free loops cut-point factorise. Accordingly, $\mathbf{P}^{1}_{x,\mathtt{C}_{n}^{d}}\cup\mathbf{P}^{1}_{x,\mathtt{C}_{n}^{d}}$ is a union of two independent cycle measures and a Bernoulli percolation and therefore, it percolates only if it has better connection probabilities in finite volume than $\phi^{1}_{x,\mathtt{C}_{n}^{d}}$. But the finite-volume two-point function of $\phi^{1}_{x,\mathtt{C}_{n}^{d}}$ is always larger than that of $\mathbf{P}^{1}_{x,\mathtt{C}_{n}^{d}}\cup\mathbf{P}^{1}_{x,\mathtt{C}_{n}^{d}}$ by (8). Since $x<x_{c}(\phi^{1}_{\mathtt{C}_{n}^{d}}),$ we conclude that $\mathbf{P}^{1}_{x,\mathtt{C}_{n}^{d}}\cup\mathbf{P}^{1}_{x,\mathtt{C}_{n}^{d}}$ does not percolate.
@@ 4. Explicit computation of critical points
[…]
@@
[…]
@@ 5. The critical probability for Bernoulli percolation is no obstruction for the UEG
The UEG of the latter model always percolates by [11, Theorem 1.3], and it is reasonable to believe that continuity of the phase transition implies that breaking even a small fraction $\delta=\delta(\varepsilon)$ of the edges breaks the infinite cluster. However, the proof we give relies on [9], the results of which are (conjecturally) not valid for all planar percolation models with a continuous phase transition (see discussion after [9, Theorem 7.5]). Therefore, the matter is much more subtle than one might expect and restating all necessary prerequisites is beyond the scope of the present paper. As such, we shall settle for referring to suitable places in the literature. However, we find the example to be important in the sense that the graph produced is, in some sense, a much more natural negative resolution to the question.
@@ 5.1. The edge-halving construction
[…]
@@
For our second example, for $x>x_{c}$ we let $\mathbb{G}^{x}$ denote the infinite cluster of $\phi_{x,\mathbb{Z}^{2}}$. By [11, Theorem 3.1], we have that the uniform even subgraph of $\mathbb{G}^{x}$ percolates almost surely. Therefore, we obtain a second proof of Theorem 1.4 if we can prove the following:
@@
It is well-known that if $\rho$ is sufficiently small, an estimate of the form (13) for some $k$ is enough to imply non-percolation by techniques that go back to [21] (see e.g. the proof of [16, Proposition 2.11]). However, by continuity of the finite volume measures, (13) remains true if $\omega_{\Lambda_{3k}}$ is replaced by $\tilde{\omega}_{\Lambda_{3k}}\sim\phi^{1}_{x_{c}+\varepsilon,\Lambda_{3k}}$ for $\varepsilon$ sufficiently small. Upon inspection, this is the same as saying that $p_{c}(\mathbb{P}_{\mathbb{G}^{x}})\geq 1-\delta$ almost surely for $x\in(x_{c},x_{c}+\varepsilon),$ which is what we wanted, since, under the increasing coupling, $x\mapsto p_{c}(\mathbb{P}_{\mathbb{G}^{x}})$ is almost surely decreasing.
Now, to see that (12) holds provided $k$ is large enough, we refer to [24, Lemma 5.2]. This lemma is stated in the context of Boolean percolation, but as is remarked on in that paper, its proof only relies on the techniques of [9, Theorem 7.5]. Thus, it is also valid for the random-cluster model. Combining [24, Lemma 5.2] with the fact that the four-arm exponent of the random-cluster model is smaller than 2 [8, Page 11] yields (12). ∎
@@ Acknowledgements
FRK acknowledge the Villum Foundation for funding through the QMATH Center of Excellence (Grant No. 10059) and the Villum Young Investigator (Grant No. 25452) programs and the Carlsberg Foundation, grant CF24-0466. UTH acknowledges funding from Swiss SNF. PW was supported by the European Research Council under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 851682 SPINRG). We thank Boris Kjær for many discussions.

### v2

@@ 2.2. Generalisations and non-uniqueness of random current phase transitions
Let $F$ be a finite graph and $v,w$ be two vertices. Let $\{\mu_{x,F}\}_{x\in[0,1]}$ be a family of cut-point factorising percolation measures. Suppose that there exists a finite graph $F$ such that $x\mapsto\mu_{x,F}[v\leftrightarrow w]$ is not monotone. Then, there exists an infinite graph $\mathbb{M}$ such that the percolation regime $\mathcal{P}(\mu_{x,\mathbb{M}})$ is disconnected.
@@
For $p\in(0,1)$ let $\mathbb{G}_{p}\sim\mathbb{P}_{p,\mathbb{T}^{d}}$. Notice that almost surely, $x_{c}(\mathbb{P}_{x,\mathbb{G}_{p}})=\frac{1}{p(d-1)}.$ By tuning the parameters $p$ and $d$ appropriately, we can make sure that
@@
Now, construct $\mathbb{M}$ by sampling $\mathbb{G}_{p}$ and substituting each edge of $\mathbb{G}_{p}$ by a copy of $F$, gluing in the same way as in our construction in the proof of Theorem 1.1. ∎
@@
As a consequence, we can see that the single random current also admits a disconnected percolation regime: Recall that [21, Figure 2.3] gives an example of a graph for which connection probabilities are not monotone, hence, by the above proposition, it follows that we get:
@@
[…]
where 0 denotes the empty graph, which is even. It is immediate that if $(G_{1},G_{2})$ is a cut-point factorisation of $G$ and $F_{1},F_{2}$ are subforests of $G_{1}$ and $G_{2}$ respectively, then $F_{1}\cup F_{2}$ is a subforest of $G$. Thus, $\nu_{\beta,G}$ cut-point factorises. At present, the connection probabilities of the model are conjectured to be monotone in $\beta$ on all finite graphs $G$ [17, p.2], but supposing that they are not, our construction would go through for this model as well, yielding a graph on which the percolation phase transition of the model is not unique.
@@ 3. Phase transitions of the wired models on the $d$-regular tree coincide
In this section, we prove the first part of Theorem 1.3 on trees. The main tool in the proof is the observation that all non-empty even subgraphs in $\mathbb{T}^{d}$ are infinite. In particular, $\ell^{1}_{x,\mathbb{T}^{d}}$ is supported on the empty graph if it does not percolate.
@@
By tail-triviality (see [16, Theorem 10.67]), $\phi^{1}_{x,\mathbb{T}^{d}}[\mathcal{C}_{\infty}]\in\{0,1\}$ and since $\phi^{1}_{x,\mathbb{T}^{d}}[\omega_{e}=0]\in(0,1),$ we conclude that
@@
For $x>x_{c}(\phi^{1}_{\mathbb{T}^{d}}),$ there exists $c>0$ such that for any vertex $v$,
@@
Consider two neighbouring vertices $v,w$ and define the event $A_{v}=\{v\longleftrightarrow\infty\text{ in }\mathbb{T}^{d}\setminus(v,w)\}$ and analogously $A_{w}$. Define the event $L_{v,w}:=A_{v}\cap A_{w}\cap\{(v,w)\text{ open}\}$. In words, this is the event that that there is a loop which contains the edge $(v,w)$ and goes through the wired boundary at infinity. Now, suppose $x>x_{c}(\phi^{1}_{\mathbb{T}^{d}})$. In this case, by the FKG inequality, we obtain a lower bound that a random-cluster configuration $\omega$ satisfies $L_{v,w}$:
@@
Second, the probability of a given edge, which is part of a loop in $\omega$, being open in $\operatorname*{UEG}_{\omega}$ is $\frac{1}{2}$ (see the much more general statement [18, Lemma 3.5]).
@@
To finish the proof of the first statement in Theorem 1.3, we make a short aside to discuss the subcritical regime of the random-cluster and random current models on the tree. It is classical that, for $x<x_{c}(\phi_{\mathbb{T}^{d}}^{1}),$ we have that $\phi_{x,\mathbb{T}^{d}}^{1}=\mathbb{P}_{x,\mathbb{T}^{d}}$ (see [16, Theorem 10.67]). A similar result holds for the double random current:
@@
Since $x<x_{c}(\phi_{\mathbb{T}^{d}}^{1})$ and $\ell^{1}_{x,\mathbb{T}^{d}}\preceq\phi^{1}_{x,\mathbb{T}^{d}}$, we get that $\ell_{x,\mathbb{T}^{d}}^{1}[\mathcal{C}_{\infty}]\leq\phi_{x,\mathbb{T}^{d}}^{1}[\mathcal{C}_{\infty}]=0.$ Therefore, since all non-singleton components of an even subgraph of $\mathbb{T}^{d}$ are infinite, we have that
@@
For $d\geq 2$, $x_{c}(\ell^{1}_{\mathbb{T}^{d}})=x_{c}(\mathbf{P}^{1}_{\mathbb{T}^{d}})=x_{c}(\mathbf{P}^{1}_{\mathbb{T}^{d}}\cup\mathbf{P}^{1}_{\mathbb{T}^{d}}).$
@@
For $a)$ if $x<x_{c}(\phi^{1}_{\mathtt{C}_{n}^{d}}),$ rather than $\eta\sim\ell^{1}_{x,\mathtt{C}_{n}^{d}}$ being deterministically empty, it includes each simple cycle of $\mathtt{C}_{n}^{d}$ independently since the free loops cut-point factorise. Accordingly, $\mathbf{P}^{1}_{x,\mathtt{C}_{n}^{d}}\cup\mathbf{P}^{1}_{x,\mathtt{C}_{n}^{d}}$ is a union of two independent cycle measures and a Bernoulli percolation and therefore, it percolates only if it has better connection probabilities in finite volume than $\phi^{1}_{x,\mathtt{C}_{n}^{d}}$. But the finite-volume two-point function of $\phi^{1}_{x,\mathtt{C}_{n}^{d}}$ is always larger than that of $\mathbf{P}^{1}_{x,\mathtt{C}_{n}^{d}}\cup\mathbf{P}^{1}_{x,\mathtt{C}_{n}^{d}}$ by (8). Since $x<x_{c}(\phi^{1}_{\mathtt{C}_{n}^{d}}),$ we conclude that $\mathbf{P}^{1}_{x,\mathtt{C}_{n}^{d}}\cup\mathbf{P}^{1}_{x,\mathtt{C}_{n}^{d}}$ does not percolate.
@@ 4. Explicit computation of critical points
$x_{c}(\mathbf{P}^{0}_{\mathtt{C}_{n}^{d}}\cup\mathbf{P}^{0}_{\mathtt{C}_{n}^{d}})=\sqrt[2n]{(2d-5)-\sqrt{(2d-5)^{2}-1} }.$
@@
$x_{c}(\mathbf{P}^{0}_{\mathtt{C}_{n}^{d}}\cup\mathbf{P}^{0}_{\mathtt{C}_{n}^{d}})=\sqrt[2n]{(2d-5)-\sqrt{ (2d-5)^{2}-1} }.$
@@ 5. The critical probability for Bernoulli percolation is no obstruction for the UEG
The UEG of the latter model always percolates by [13, Theorem 1.3], and it is reasonable to believe that continuity of the phase transition implies that breaking even a small fraction $\delta=\delta(\varepsilon)$ of the edges breaks the infinite cluster. However, the proof we give relies on [10], the results of which are (conjecturally) not valid for all planar percolation models with a continuous phase transition (see discussion after [10, Theorem 7.5]). Therefore, the matter is much more subtle than one might expect and restating all necessary prerequisites is beyond the scope of the present paper. As such, we shall settle for referring to suitable places in the literature. However, we find the example to be important in the sense that the graph produced is, in some sense, a much more natural negative resolution to the question.
@@ 5.1. The edge-halving construction
$\smash{\overset{\eta}{v\longleftrightarrow w}\quad\text{if and only if}\quad\overset{\psi(\eta)}{v\longleftrightarrow w}}$,
@@
For our second example, for $x>x_{c}$ we let $\mathbb{G}^{x}$ denote the infinite cluster of $\phi_{x,\mathbb{Z}^{2}}$. By [13, Theorem 3.1], we have that the uniform even subgraph of $\mathbb{G}^{x}$ percolates almost surely. Therefore, we obtain a second proof of Theorem 1.4 if we can prove the following:
@@
It is well-known that if $\rho$ is sufficiently small, an estimate of the form (13) for some $k$ is enough to imply non-percolation by techniques that go back to [23] (see e.g. the proof of [18, Proposition 2.11]). However, by continuity of the finite volume measures, (13) remains true if $\omega_{\Lambda_{3k}}$ is replaced by $\tilde{\omega}_{\Lambda_{3k}}\sim\phi^{1}_{x_{c}+\varepsilon,\Lambda_{3k}}$ for $\varepsilon$ sufficiently small. Upon inspection, this is the same as saying that $p_{c}(\mathbb{P}_{\mathbb{G}^{x}})\geq 1-\delta$ almost surely for $x\in(x_{c},x_{c}+\varepsilon),$ which is what we wanted, since, under the increasing coupling, $x\mapsto p_{c}(\mathbb{P}_{\mathbb{G}^{x}})$ is almost surely decreasing.
Now, to see that (12) holds provided $k$ is large enough, we refer to [27, Lemma 5.2]. This lemma is stated in the context of Boolean percolation, but as is remarked on in that paper, its proof only relies on the techniques of [10, Theorem 7.5]. Thus, it is also valid for the random-cluster model. Combining [27, Lemma 5.2] with the fact that the four-arm exponent of the random-cluster model is smaller than 2 [9, Page 11] yields (12). ∎
@@ Acknowledgements
FRK acknowledge the Villum Foundation for funding through the QMATH Center of Excellence (Grant No. 10059) and the Villum Young Investigator (Grant No. 25452) programs and the Carlsberg Foundation, grant CF24-0466. UTH acknowledges funding from Swiss SNF. PW was supported by the European Research Council under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 851682 SPINRG). We thank Boris Kjær and Franco Severo for many discussions and comments.